OpenStudy (anonymous):

Find the sum of a 9-term geometric sequence when the first term is 4 and the last term is 1,024 and select the correct answer below. 682 2,044 2,048 678

OpenStudy (campbell_st):

well you need to find the common ratio... you know the 9th term is 1024 and the 1st term is 4 so use the formula for a term \[a_{n} = a_{1} \times r^{n - 1}\] so \[1024 = 4\times r^8\] solve for r. the fund the sum of 9 terms using \[S_{9} = \frac{4( 1 - r^9)}{1 - r}\] hope it helps